# Learning to Control Unknown Strongly Monotone Games

### v2

@@ Abstract
Consider a game where the players’ utility functions include a reward function and a linear term for each dimension, with coefficients that are controlled by the manager. We assume that the game is strongly monotone, so gradient play converges to a unique Nash equilibrium (NE). The NE is typically globally inefficient. The global performance at NE can be improved by imposing linear constraints on the NE. We therefore want the manager to pick the controlled coefficients that impose the desired constraint on the NE. However, this requires knowing the players’ reward functions and action sets. Obtaining this game information is infeasible in a large-scale network and violates user privacy. To overcome this, we propose a simple algorithm that learns to shift the NE to meet the linear constraints by adjusting the controlled coefficients online. Our algorithm only requires the linear constraints violation as feedback and does not need to know the reward functions or the action sets. We prove that our algorithm converges with probability 1 to the set of NE that satisfy target linear constraints. We then prove an L2 convergence rate of near-$O(t^{-1/4})$.
@@ 1 Introduction
Large-scale networks such as in communication, transportation, energy, and computing consist of myriad of agents taking local decisions. Agents do not see the bigger picture therefore optimize their local objectives instead. The result is often an inefficient equilibrium. Hypothetically, a centralized authority that knows all network parameters can instruct the agents how to act to optimize a global objective. However, such a centralized architecture is infeasible for large-scale networks and also violates the users’ privacy. As a result, there has been a surge of interest in distributed multiagent algorithms [1, 2, 3, 4, 5].
Game theory is a natural way to tackle distributed decision-making [6, 7]. The game structure formalizes the effect of the interaction between users. The players may be cooperative, or they may be selfishly interested in maximizing their reward. In large-scale systems, obtaining performance guarantees is challenging even with cooperative players since they do not know the reward functions of their peers. A Nash Equilibrium (NE) aims to predict the outcome of such a repeated interaction. In monotone games [8, 9, 10], which are our focus here, simple distributed algorithms such as gradient descent of players on their reward functions are guaranteed to converge to a NE.
A NE does not optimize global objectives and can lead to poor performance. By controlling parameters in the utility functions of the players, the manager can aim to change the game such that the NE is efficient. In general, setting utility parameters to optimize performance at the NE is intractable. Therefore, in this paper, we focus on the task of imposing linear constraints on the resulting NE of the game the manager controls. Linear constraint are common in many optimization formulations. When optimizing a global quadratic cost, linear constraint can describe the first-order optimality condition for the NE. In resource allocation, linear constraints can describe what should be the ideal loads at each resource at NE.
@@
To converge to a NE that meets target linear constraints in an unknown strongly monotone game, we propose an online learning approach. In our algorithm, the manager adjusts the controlled coefficients online and receives the constraint violation as feedback. Such feedback is easy to monitor in practice and maintains user privacy. Our algorithm is based on two time-scale stochastic approximation, where the agents update their actions in the faster time-scale and the manager updates the controlled coefficients in the slower time-scale.
We prove that our algorithm converges to the set of NE that satisfy the target linear constraints with probability 1. We further show that our algorithm has a mean square error rate of near-$O(t^{-1/4})$ for the linear constraint. Due to the convex and compact action sets, our slower time-scale is a fixed point iteration with a non-expansive mapping instead of the stronger contractive mapping typically assumed to analyze convergence rates [11, 12]. Apart from stochastic approximation, we employ tools from variational inequalities and the Krasnosel’skiĭ–Mann (KM) algorithm in our analysis. Therefore, we provide a theoretically principled game control scheme for strongly monotone games.
@@ 1.1 Related Work
Distributed optimization [13, 14] is concerned with optimizing a target objective distributedly over a network of agents. However, these agents are non-interacting computational units so there is no game structure between them. To cast the game as a distributed optimization with the action profile as the variable, each agent would need to know the effect of other agents on its reward function, which is infeasible. Moreover, distributed optimization entails significant communication overhead between agents.
Distributed algorithms that converge to NE have been studied in [15, 16, 17], including in monotone games [18, 19, 20]. The players in our game use simple stochastic gradient play to converge to NE. However, if the game is uncontrolled, the NE is typically globally inefficient. Using our game control scheme, the manager learns to steer the game into an efficient NE (i.e., a NE that satisfies target linear constraints).
Compared to mechanism design approaches [21, 22], we focus on cooperative or myopic players and our manager does not elicit private information from the players but only observes the constraint violation.
@@
Our learning framework can be viewed as a bandit learning problem where the manager is playing against a bandit that is not stochastic nor fully adversarial [26, 27]. Instead, the rewards of this bandit are generated by the game dynamics. This “game bandit” introduces a new type of noise. The goal of the manager is to steer the unique NE to a point that satisifes target linear constraints. Hence, the type of feedback the manager needs concerns the behavior at NE. However, the dynamics are not at NE every turn. In fact, by changing the prices the manager perturbs the system which complicates and delays the convergence to the unique NE. Therefore, the main noise in this “game bandit” is an equilibrium noise that stems from the distance of the dynamics from the time-varying NE.
Our problem can be viewed as a Generalized Nash Equilibrium Problem (GNEP), where players seek to converge to a NE subject to coupled constraints [28]. Efficient computation of GNE in strongly monotone games in the noiseless setting has been widely studied [29, 30, 31]. Stochastic convergence to GNEs has recently been studied under different assumptions. In [32] players share their actions with others, and in [33] the reward function is assumed to be linear. In [34], stochastic convergence to GNE is studied more generally, resulting in a centralized algorithm where the weighted empirical average of the actions is shown to converge in expectation. We provide a distributed stochastic learning scheme that converges in probability 1 to GNE with linear coupled constraints in strongly monotone games without communication between players.
Our work belongs to the literature on intervention and control of games [35, 36, 37, 38, 39, 40, 41, 42, 43]. In [44] it was assumed that the manager knows the reward functions of the players, and can compute the desired NE. This is also the assumption in [45, 39], that considered the general approach of game-theoretic control. In [35], an aggregative game where the cost is a linear function of the aggregated action was considered. It was assumed that the manager can control the total loads the players observe, which can deviate from the true total loads. In [46, 47, 48], a similar approach to ours was considered for discrete congestion games, where the reward function of all players is the same convex function of the number of players that share the chosen resource. The congestion game considered there is a special case of monotone games that allow us to model far more general interactions and action sets. For example, it allows for models when not only the number of players that share a resource matters but also their identity (e.g., geometric locations). Additionally, the pricing scheme in [46, 47, 48] requires the manager to know the reward functions. Control of dynamic Markov potential games was considered in [40], where it was assumed that the manager knows the players’ reward functions and can simulate the game.
@@
An earlier conference paper [50] was limited to load balancing in resource allocation. This paper’s framework is similar to that in [50], but is extended to linear constraint in strongly monotone games. The NE guarantees here are of the form $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$ instead of just $\sum_{n=1}^{N}\boldsymbol{x}_{n}\left(\boldsymbol{\alpha}^{*}\right)=\boldsymbol{\ell^{*}}$. Nevertheless, the results here are completely novel and independent of [50]. Our convergence guarantees are with probability 1 instead of in L2 like in [50]. Unlike [50], we provide a finite time convergence error bound. Our novel analysis is based on two time-scale stochastic approximation with non-expansive mappings.
@@ 1.2 Outline and Notation
The rest of the paper is organized as follows. Section 2 formulates our game control problem and details our assumptions. Subsection 2.1 details two applications that fall under our general model. Section 3 describes the game control algorithm and states our main results regarding its performance guarantees. Section 4 breaks the proofs of our main results into lemmas. Section 5 provides numerical simulations for the application described earlier. Finally, Section 6 concludes the paper and proposes future research directions.
[…]
@@ 2 Problem Formulation
$u_{n}(\boldsymbol{x};\boldsymbol{\beta})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{i}x_{n}^{i}.$ (1)
Here $r_{n}(\boldsymbol{x})$ is the reward function for player $n$. We assume that $r_{n}(\boldsymbol{x})$ is twice continuously differentiable for each $n$ and is greater than or equal to 00 for all $\boldsymbol{x}\in\mathcal{X}$. The vector $\boldsymbol{\beta}\in\mathbb{R}^{Nd}$ denotes the ‘controlled coefficients’ controlled by the game manager. We use the same vector notation for $\boldsymbol{\beta}$ and $\boldsymbol{\beta}_{n}$.
@@
The key assumption about the game structure is that it is a strongly monotone game. For this, we define the gradient operator $F(\cdot):\mathbb{R}^{Nd}\mapsto\mathbb{R}^{Nd}$ as
@@
There exists $\mu>0$ such that for all $\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
@@
The above assumption means that the game is a strongly monotone game with parameter $\mu>0$. Monotone games are a well-studied class of games [51]. The linear term in (1) does not affect the strongly monotone property of the game since the gradient vector becomes $F(\boldsymbol{x})-\boldsymbol{\beta}$. Since it is a strongly monotone game, there exists a unique pure NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$ corresponding to each vector $\boldsymbol{\beta}$, defined as:
@@
A strategy profile $(\boldsymbol{x^{*}}_{n}(\boldsymbol{\beta}),\boldsymbol{x^{*}}_{-n}(\boldsymbol{\beta}))$ is a pure Nash equilibrium (NE) corresponding to coefficients $\boldsymbol{\beta}$, if $u_{n}(\boldsymbol{x}_{n}^{\boldsymbol{*}}(\boldsymbol{\beta}),\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})\geq u_{n}(\boldsymbol{x}_{n},\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})$, for all $\boldsymbol{x}_{n}\in\mathcal{X}$ and all $n\in\mathcal{N}$.
We make the following additional assumption about the game, which says that the gradient function is Lipschitz continuous. It is merely a technical assumption, commonly taken along with the strongly monotone condition [52, 53].
@@
There exists $L>0$ such that for all $\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
@@
Our game evolves in discrete time, so at time $t$ players take actions $\boldsymbol{x}_{n,t}$ given the controlled coefficients $\boldsymbol{\beta}_{n,t}$. The players update their actions using gradient ascent and the controlled coefficients are adjusted based on the manager’s objective (discussed below (6)). Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., uncontrolled game), the players will converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$.
In many scenarios, the players do not fully know their reward functions but instead learn them on the fly based on the reward values they receive (i.e., bandit feedback) or other data they collect. The gradient can then be approximated from this data using regression or other supervised machine learning techniques [54]. These learning schemes make the players’ actions stochastic which adds another type of noise to our analysis. We therefore assume that agents observe noisy variants of their gradients, where the noise is a martingale difference sequence with a bounded second moment.
@@
Each player $n$ observes $\boldsymbol{g}_{n,t-1}$ at time $t$, where $\boldsymbol{g}_{n,t-1}=\nabla_{\boldsymbol{x}_{n,t-1}}r_{n}(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{n,t}$. Then $\boldsymbol{g}_{t-1}=F(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{t}$, where $\boldsymbol{M}_{t}$ satisfies

### v3

@@ Abstract
Consider a strongly monotone game where the players’ utility functions include a reward function and a linear term for each dimension, with coefficients that are controlled by the manager. Gradient play converges to a unique Nash equilibrium (NE) that does not optimize the global objective. The global performance at NE can be improved by imposing linear constraints on the NE, also known as a generalized Nash equilibrium (GNE). We therefore want the manager to control the coefficients such that they impose the desired constraint on the NE. However, this requires knowing the players’ rewards and action sets. Obtaining this game information is infeasible in a large-scale network and violates user privacy. To overcome this, we propose a simple algorithm that learns to shift the NE to meet the linear constraints by adjusting the controlled coefficients online. Our algorithm only requires the linear constraints violation as feedback and does not need to know the reward functions or the action sets. We prove that our algorithm converges with probability 1 to the set of GNE given by coupled linear constraints. We then prove an L2 convergence rate of near-$O(t^{-1/4})$.
@@ 1 Introduction
Large-scale networks such as in communication, transportation, energy, and computing consist of a myriad of agents taking local decisions. Agents do not see the bigger picture and instead only optimize their local objectives. The result is often an inefficient equilibrium [1]. Hypothetically, a centralized authority that knows all network parameters can instruct the agents how to act to optimize a global objective. However, such a centralized architecture is infeasible for large-scale networks and also violates the users’ privacy.
The interaction between the users is formalized as a game [2, 3]. The players may be cooperative, or they may be selfishly interested in maximizing their reward. In large-scale systems, obtaining performance guarantees is challenging even with cooperative players since they do not know the reward functions of their peers. A Nash Equilibrium (NE) aims to predict the outcome of such a repeated interaction. In monotone games [4, 5, 6], which are our focus here, simple distributed algorithms such as gradient ascent of players on their reward functions are guaranteed to converge to a NE.
A NE does not optimize global objectives and can lead to poor performance. By controlling parameters in the utility functions of the players, the manager can aim to change the game such that the NE is efficient. In general, setting utility parameters to optimize performance at the NE is intractable. Therefore, in this paper, we focus on the task of imposing linear constraints on the resulting NE of the game the manager controls. Linear constraints are common in many optimization formulations. When optimizing a global quadratic cost, linear constraints can describe the first-order optimality condition for the NE. In resource allocation, linear constraints can describe what should be the ideal loads at each resource at NE.
@@
To converge to a NE that meets target linear constraints in an unknown strongly monotone game, we propose an online learning approach. In our algorithm, the manager adjusts the controlled coefficients online and receives the constraint violation as feedback. Such feedback is easy to monitor in practice and preserves user privacy. Our algorithm is based on two time-scale stochastic approximation, where the agents update their actions in the faster time-scale and the manager updates the controlled coefficients in the slower time-scale.
We prove that our algorithm converges to the set of NE that satisfy the target linear constraints with probability 1. We further show that our algorithm has a mean square error rate of near-$O(t^{-1/4})$ for the linear constraints. Due to the convex and compact action sets, our slower time-scale is a fixed point iteration with a non-expansive mapping instead of the stronger contractive mapping that allows for a $O(t^{-1})$ convergence rate [7, 8]. Apart from stochastic approximation, we employ tools from variational inequalities and the Krasnosel’skiĭ–Mann (KM) algorithm in our analysis. Therefore, we provide a theoretically principled game control scheme for strongly monotone games.
@@ 1.1 Related Work
Distributed optimization [9, 10] is concerned with optimizing a target objective over a network of agents. However, these agents are non-interacting computational units that are not coupled in a game structure. To cast the game as a distributed optimization with the action profile as the variable, each agent would need to know the effect of other agents on its reward, which is infeasible. Moreover, distributed optimization requires significant communication between agents.
Distributed algorithms that converge to NE have been studied in [11, 12, 13], and in [14, 6, 15] for monotone games. The players in our game use simple stochastic gradient play to converge to NE. However, if the game is uncontrolled, the NE does not optimize global objectives. In our game control scheme, the manager learns to steer the game to an efficient NE, which is a GNE with target linear constraints.
Convergence to GNE [16] in strongly monotone games is usually studied in the noiseless setting [17, 18, 19], while our learning agents make the convergence to a GNE stochastic. Stochastic convergence to GNEs has recently started to attract attention. In [20] players share their actions with others, and in [21] the reward function is assumed to be linear. Closer to our scenario is [22], which studied a centralized algorithm in a monotone game where the weighted empirical average of the actions is shown to converge in expectation to a GNE. In contrast, our GNE is limited to linear coupled constraints in strongly monotone games. For this scenario, we provide a distributed stochastic learning scheme that converges in probability 1 to a GNE, without communication between players. To the best of our knowledge, our work is the first to study distributed and stochastic convergence to a GNE.
Our algorithm is a stochastic iterative approach to solving a constrained optimization problem using Lagrange multipliers. Convergence guarantees for stochastic augmented Lagrangian schemes are scarce. In [23], the continuous-time optimization problem was studied, and [24] provides asymptotic results for a variant of our algorithm. We present the convergence rate for our algorithm using novel analysis techniques.
Unlike mechanism design [25, 26], our players are cooperative or myopic, and our manager does not elicit any players’ private information, but only observes the constraint violation.
@@
Our manager is learning with bandit feedback that is not stochastic nor fully adversarial [30, 31]. Instead, the rewards of this bandit are generated by the game dynamics. This “game bandit” introduces a new type of noise. The goal of the manager is to steer the unique NE to a point that satisfies target linear constraints. Hence, the feedback the manager needs concerns the behavior at NE. However, the dynamics are not at NE every turn. In fact, by changing the prices, the manager perturbs the system which complicates and delays the convergence to the unique NE. Therefore, constraint violation feedback in this problem suffers from equilibrium noise which is the distance of the dynamics from the time-varying NE.
Our work belongs to the literature on intervention and control of games [32, 33, 34, 35, 36, 37, 38, 39, 40]. In [41] it was assumed that the manager knows the reward functions of the players, and can compute the desired NE. This is also the assumption in [42, 36], that considered the general approach of game-theoretic control. In [32], an aggregative game where the cost is a linear function of the aggregated action was considered. It was assumed that the manager can control the total loads the players observe, which can deviate from the true total loads. In [43, 44, 45], a similar approach to ours was considered for discrete congestion games, where the cost function of all players is the same convex function of the number of players that share the chosen resource. The congestion game considered there is a special case of monotone games that allow us to model far more general interactions and action sets. For example, it allows for models when not only the number of players that share a resource matters but also their identity (e.g., geometric locations). Additionally, the pricing scheme in [43, 44, 45] requires the manager to know the reward functions. Control of dynamic Markov potential games was considered in [37], where it was assumed that the manager knows the players’ reward functions and can simulate the game.
@@
Our convergence rate is limited by the non-expansive mapping in the slower time-scale, and not by the algorithm players use to update their actions. This non-expansive mapping arises due to the projection of the actions into the action sets in the faster time-scale. When the action sets are $\mathbb{R}^{d}$, our algorithm becomes a two-time-scale iteration with a contractive mapping in both time-scales, with a convergence rate of $O(t^{-1})$ [7]. This slowdown is a known challenge in the literature. While the optimal bound known for single time-scale fixed point iterations with contractive mappings is $O(t^{-1})$ [8], the best known bound for iterations with non-expansive mappings is $O(t^{-1/3})$ [47]. Compared to $O(t^{-1/3})$, we attribute the slightly slower convergence rate to the two-time scale dynamics.
An earlier conference paper [48] was limited to load balancing in resource allocation. This paper extends [48] to linear constraints in strongly monotone games. The NE guarantees here are of the form $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$ instead of just $\sum_{n=1}^{N}\boldsymbol{x}_{n}\left(\boldsymbol{\alpha^{*}}\right)=\boldsymbol{\ell^{*}}$. Nevertheless, the results here are novel and independent of [48]. Our convergence guarantees hold both almost surely and in $L_{2}$, whereas [48] establishes only $L_{2}$ convergence. Unlike [48], we provide a finite time convergence error bound. Our novel analysis uses two time-scale stochastic approximation with non-expansive mappings.
@@ 1.2 Outline and Notation
The rest of the paper is organized as follows. Section 2 formulates our game control problem and details our assumptions. Subsection 2.1 details two applications that fall under our general model. Section 3 describes the game control algorithm and states our main results regarding its performance guarantees. Section 4 breaks the proofs of our main results into lemmas. Section 5 provides numerical simulations for the applications described earlier. Finally, Section 6 concludes the paper and proposes future research directions.
We use bold letters to denote vectors and matrices, and the standard game-theoretic notation where $\boldsymbol{x}_{-n}$ is the vector of actions for all players except player $n$. $\|\cdot\|$ denotes the $\ell_{2}$-norm throughout the paper when applied on a vector, and the operator norm for matrices (under the $\ell_{2}$-norm for vectors).
@@ 2 Problem Formulation
$u_{n}(\boldsymbol{x};\beta_{n}^{1},\ldots,\beta_{n}^{d})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{i}x_{n}^{i}.$ (1)
Here $r_{n}(\boldsymbol{x})$ is the reward function for player $n$. We assume that $r_{n}(\boldsymbol{x})$ is twice continuously differentiable for each $n$. The vector $\boldsymbol{\beta}\in\mathbb{R}^{Nd}$ denotes the ‘controlled coefficients’ controlled by the game manager. We define the same vector notation for $\boldsymbol{\beta}$ and $\boldsymbol{\beta}_{n}$ as we do for $\boldsymbol{x}$ and $\boldsymbol{x}_{n}$, respectively. Such a linear term arises when introducing price per use of a resource, or when agents are cooperative and can compute such a term since they know their own action.
@@
The key assumption is that the game is strongly monotone. For this, define the gradient operator $F(\cdot):\mathbb{R}^{Nd}\mapsto\mathbb{R}^{Nd}$:
@@
There exists $\mu>0$ such that $\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
@@
Thus, the game is strongly monotone with parameter $\mu>0$, or equivalently the function $F(\cdot)$ is $\mu$-strongly monotone. Monotone games are a well-studied class of games [49]. The linear term in (1) does not affect the strongly monotone property of the game since the gradient vector becomes $F(\boldsymbol{x})-\boldsymbol{\beta}$. In this strongly monotone game, there exists a unique pure NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$ corresponding to $\boldsymbol{\beta}$, defined as:
@@
An action profile $(\boldsymbol{x^{*}}_{n}(\boldsymbol{\beta}),\boldsymbol{x^{*}}_{-n}(\boldsymbol{\beta}))$ is a pure Nash equilibrium (NE) corresponding to coefficients $\boldsymbol{\beta}$, if $u_{n}(\boldsymbol{x}_{n}^{\boldsymbol{*}}(\boldsymbol{\beta}),\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})\geq u_{n}(\boldsymbol{x}_{n},\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})$, for all $\boldsymbol{x}_{n}\in\mathcal{X}$ and all $n\in\mathcal{N}$.
We assume that the gradients of the players are Lipschitz continuous, which is a standard assumption [50, 51].
@@
There exists $L>0$ such that $\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
@@
Our game evolves in discrete time, so at time $t$ player $n$ plays $\boldsymbol{x}_{n,t}$ given the controlled coefficients $\boldsymbol{\beta}_{n,t}$. The players update their actions using gradient ascent and the controlled coefficients are adjusted based on the manager’s objective (discussed below (6)). Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., uncontrolled game), the players will converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$ [4].
In many scenarios, players do not fully know their reward functions but instead learn them on the fly based on the reward values they receive or other data they collect. The gradient can then be approximated from this data [52]. These learning schemes make the players’ actions stochastic, which adds another type of noise to our analysis.
@@
Each player $n$ observes $\boldsymbol{g}_{n,t-1}$ at time $t$, where $\boldsymbol{g}_{n,t-1}=\nabla_{\boldsymbol{x}_{n}}r_{n}(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{n,t}$. Then $\boldsymbol{g}_{t-1}=F(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{t}$, where $\boldsymbol{M}_{t}$ satisfies
@@
The idea behind this assumption is that from the point of view of the players, the rewards are stochastic, i.e., $r_{n}(\boldsymbol{x})\triangleq r_{n}(\boldsymbol{x};\xi)$ for an unknown random variable $\xi$. Then the optimal estimator of the gradient of player $n$ from past observations is $\boldsymbol{g}_{n,t}\triangleq\mathbb{E}\left\{\nabla_{\boldsymbol{x}_{n}}r_{n}(\boldsymbol{x}_{t})\mid\mathcal{F}_{t}\right\}$. This makes the estimation error $\boldsymbol{M}_{n,t+1}=\boldsymbol{g}_{n,t}-\nabla_{\boldsymbol{x}_{n}}r_{n}(\boldsymbol{x}_{t})$ a Martingale difference sequence, and $\left\|\boldsymbol{M}_{n,t}\right\|$ is bounded if $\nabla_{\boldsymbol{x}_{n}}r_{n}(\boldsymbol{x})$ is bounded. Gradient estimation from bandit feedback does not satisfy Assumption 3, which requires unbiased estimates.